A VCR and TV were bought for ₹ 8,000 each. The shopkeeper made a loss of 4% on the VCR and a profit of 8% on the TV. Find the gain or loss percent on the whole transaction. 

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

C.P. of a VCR =₹8000
The shopkeeper made a loss of 4 % on VCR.
This means if C.P. is₹100, then S.P. is₹96.
When C.P. is₹8000,
S.P. =(96/100 × 8000)=₹7680
C.P.ofaTV=₹8000
The shopkeeper made a profit of 8 % on TV.
This means that if C.P. is₹100, then S.P. is₹108.
When C.P. is₹8000,
S.P. =(108/100 × 8000)=₹8640
Total S.P. =₹7680 +₹8640 =₹16320
Total C.P. =₹8000 +₹8000 =₹16000
Since, total S.P.> total C.P⟹profit.
Profit =₹ 16320 − ₹16000 =₹320
∴ The shopkeeper had a gain of 2% on the whole transaction.

Thanks